A hole is made on the wall of cylindrical vessel, containing a liquid upto a height h. Water is coming out through the hole and falls at a distance ‘t’ from the base of the cylinder. t is found to be maximum, when hole is created at depth of 2xh. Find the value of x?


1
Expert's answer
2021-02-11T17:09:42-0500

Horizontal velocity of water out of hole,


"u=\sqrt{2g(2xh)}=2\sqrt{gxh}"

Height of hole from ground level "h-2xh". The time taken buy water to cover vertical distance will be


"(1-2x)h=0.5gs^2\\s=\sqrt{\frac{2(1-2x)h}{g}}"

Range is


"t=2\sqrt{gxh}\sqrt{\frac{2(1-2x)h}{g}}=2h\sqrt{2x(1-2x)}"

Horizontal range will be maximum if


"x=\frac{1}{4}"